A ray of light is incident in water on a plane water-glass boundary. The angle of incidence is 30°.

   Calculate the angle of deviation.


1
Expert's answer
2021-03-09T07:43:01-0500

Refractive index of water wrt air = 4/3

Refractive index of glass wrt air = 3/2.

Refractive index of glass wrt water = 3/2 ×3/4 = 9/8.

Refractive index = (sin i)/(sin r) =9/8

(sin 30)/(sin r) = sin r = (8/9)sin 30


(4/9) = r = arcsin (4/9) =26.39 degrees.


Therefore

Angle of deviation is = 30 - 26.39 = 3.61degrees